Express 1 + Ö3i in exponential form
Thank you in advance for your help

Answer by venugopalramana(3286)

Express 1 + Ö3i in exponential form
WHAT IS THIS O?..I AM TAKING IT AS 1+3i ONLY
LET Z=1+3i
|Z|=SQRT(1^2+3^2)=SQRT(10)=R
ARGUMENT OF Z = TAN ^-1(3/1)=TAN^-1(3)=1.24 RADIANS ..THEN
Z=R*E^(1.24i)=SQRT(10)*E^(1.24i)

Thank you in advance for your help